OpenStudy (anonymous):

Need help with system: -7x-8y=9 -4x+9y=-22

OpenStudy (anonymous):

multiply the bottom equation by -2 and move it up 8x - 18y = 44 -7x -8y = 9 8 -18 44 -7 -8 9 add row 2 to row 1 1 -26 53 -7 -8 9 add 7 times row 1 to row 2 1 -26 53 0 -190 380 divide row 2 by -190 1 -26 53 0 1 -2 add 26 times row 2 to row 1 1 0 1 0 1 -2 x = 1 and y = -2

OpenStudy (anonymous):

thanks